1. Simplify the rational expression: Please show work
x2 – 6x – 7
x2 – 1


2. Divide:
x2 – 3x + 2÷x2 – 4
8x – 8 5x + 10


3. Simplify:
√[100x2y16z8]

4. Perform the indicated operations:
2√[72] + √[32] – 4√[18]

5. Multiply:
(√[3] + 4√[5])(2√[3] – √[5])

6. Rationalize the denominator:
3____
√[6] – √[3]

Answer by checkley71(8403)

x^2-6x-7=0
(x-7)(x+1)=0
x-7=0
x=7 solution
x+1=0
x=-1 solution
----------------------------------
x^2-1=0
(x-1)(x+1)
x-1=0
x=1 solution
x+1=0
x=-1 solution
-------------------------------------------------------
MERRY CHRISTMAS